what is the perimiter of a 30 60 90 triangle with a hypotenuse of 18
Answer by lwsshak3(11628)

what is the perimeter of a 30 60 90 triangle with a hypotenuse of 18
In a right triangle the side opposite the 30º angle is always one-half the hypotenuse=18/2=9
Use pythagorean theorem to find the third side=x
x^2=18^2-9^2=243
x=√243=15.6..
Perimeter=18+9+15.6=42.6